MULTIPLE ELASTIC CONFIGURATIONS ASSOCIATED WITH LOCALIZED VIBRATIONAL MODES IN CRYSTALS

\underline{A. J. SIEVERS}, Laboratory of Atomic and Solid State Physics, Cornell University, Ithaca, NY 14853-2501.

By examining the far infrared properties of lattice vibrations associated with defects in crystals a wide variety of spectroscopic signatures have been identified, ranging from localized gap modes and resonant modes to tunneling states. The recent discovery of multiple elastic configurations for some defect-lattice combinations points to another characteristic property of these large dynamical systems. Work supported by NSF-DMR and NASA.